# Grounding Synthetic Data Evaluations of Language Models in Unsupervised Document Corpora

### v2

@@ Appendix A Expanded related works
When constructing knowledge or fact based evaluation benchmarks, there are two broad categories of answers: multiple choice and open ended. Open-ended questions provide the model less information and require synthesis [22], but at the cost of needing another LM to evaluate the correctness of the results (LM-as-a-Judge [18]) [23]. Multiple choice question (MCQ) results can be parsed and graded more simply. Modern evaluation frameworks like Inspect-AI [26] enable shuffling answer options and contain default prompts for both open-ended and MCQ evaluations. With MCQ, having the answers visible to the LM can artificially inflate performance [24], even going as far as the model identifying the correct answers without seeing the question [25]. Both open-ended and multiple choice questions have weaknesses to question reformatting and phrasing where semantically identical questions produce different answers [27]. The simplest form of open-ended evaluation relies on a judge LM, though other frameworks have been proposed which attempt to improve the precision of open-ended evaluations [28, 29].
Appendix B Impact of question and answer length
@@
While longer more detailed questions (both MC and OE) produce a model over-performance (see Figure 5), the same trend does not appear in answer length. Figure 7 demonstrates this via scatterplot showing synthetic benchmark accuracy over-performance. Only the gpt-4.1-mini at the very right of the plot is significantly biased above the expected trend line. The over-performance trend shown for questions in Figure 5(b) is absent.
@@
The model gpt-4.1-mini demonstrates the largest over-performance for both question length and answer length as shown in Figure 5(b). The following question answer pair was generated by gpt-4.1-mini (37 tokens) using a context from Squadv2, highlighting the difference in model verbosity in both question and answers:
@@
Appendix C Synthetic benchmark question diversity
@@
Most LMs produce similar diversity scores that are closely aligned with the human questions Some highly capable models (such as gpt-4.1) show comparatively worse question diversity while performing well in clarity, correctness, and groundedness. Figure 8 shows the question diversity per generating model along with the original question diversity.
Appendix D Multiple-choice vs open-ended questions
@@
The generating models which create overly easy questions by including too much detail (like phi-4 and gpt-4.1-mini) do so for both open ended and multiple choice questions. The long question problem becomes even more pronounced for open ended questions as shown in Figure 9.
@@
These overly long synthetic questions result in an over-performance by smaller models on the generated benchmark as shown in Figure 10. We hypothesize this happens because the additional detail in the question provides the LM under evaluation more information to pattern match against text seen in the training phase, thereby increasing the likelihood that the LM retrieves memorized patterns relevant to the question. This effect is seen in Figure 10(a) for multiple choice, and Figure 10(b) for open-ended. Both plots in Figure 10 demonstrate LM benchmark evaluation accuracy over-performance ().
@@
While similar trends in question length and the over-performance that induces, the evaluation accuracy correlations and model rank order work equally well for open ended and multiple choice questions. Figure 11(a) plots the correlation in evaluation accuracy between the human reformatted questions and the synthetic generated benchmark for multiple choice, and Figure 11(b) for open-ended. Figure 11(a) shows a sec_qa() dataset specific oddity: the over-performance on human questions (top right) stems from the multiple choice human questions being very easy and produces an approximately flat line over various synthetic versions generated by LMs. Interestingly enough, this over-performance disappears under the open-ended question regime for this dataset as seen in Figure 11(b).
Appendix E Evaluation benchmark accuracy and ranking correlation
@@
Figure 11(a) shows the evaluation accuracy scatterplot comparing the reformatted human questions against the synthetic generated benchmark questions for the 5 bolded ensemble models in Table 2. This ensemble produces a Pearson accuracy correlation of between the evaluation accuracies for synthetic and human benchmarks. Each plot marker is the average evaluation accuracy for a given model-dataset combination. While this model ensemble produces reasonable overall accuracy correlation, occasionally specific LMs generate undesirable benchmarks. Figure 12 showcases a failure where the synthetic benchmark generated by gpt-4.1-mini contains overly easy questions where all models perform near evaluation accuracy (Pearson accuracy correlation .). This over-performance is likely linked to the question length problem.
@@
Figure 13 plots the evaluation model performance across all datasets, comparing the synthetic and human accuracy Spearman ranking correlation of the evaluation model capability. This matches the use case of a user selecting between various models for a specific application relevant to the provided grounding documents. Figure 13(a) is the rank order barchart equivalent of Figure 11(a) and Figure 13(b) is the rank order barchart equivalent of the failure case in Figure 12.
@@
Table 4 showcases the Spearman rank correlation between evaluation benchmarks written by humans and the synthetic data versions. The table includes data for all three of the human question reformatting models used to demonstrate the minimal impact that reformatting has on the overall high level results. Overall, the best outcome is to ensemble together a few questions generated by various reasonably capable LMs to avoid the known slight self-bias/self-preference issue. While we used Llama-4-Maverick as the text feature extractor for measures like groundedness and clarity, that model was unexpectedly dominant in these final results.
@@
Table 5 contains the benchmark evaluation Pearson accuracy correlation between the reformatted human questions and the synthetic data generated questions. Similarly to the Spearman ranking, Llama-4-Maverick performed very well in these tests. However, unlike the ranking data, using a ensemble of models produces significantly worse Pearson accuracy correlation numbers under some circumstances. Likely due to the long and overly easy questions written by gpt-4.1 and o4-mini.
@@
Appendix F Question clarity and groundedness
@@
Appendix G Case study on two academic papers
Our approach was demonstrated on two academic papers: “Current Solutions and Future Trends for Robotic Prosthetic Hands” (labeled annurev-control-071020-104336) [44] and “Recent Advances in Large Language Model Benchmarks against Data Contamination: From Static to Dynamic Evaluation” (labeled arXiv_2502_17521v1) [45]. Figure 14 showcases LM benchmark evaluation performance on these two datasets; the MCQ results show most models performing well before a sharp dropoff with 4B and fewer parameter models and the Open-Ended results with a much smoother decline in performance. Multiple-choice questions can be easier to answer due to the presence of the answer options, even if the evaluation framework shuffles the option order. This effect is likely the cause of the MCQ evaluation accuracy to remaining higher for longer.
The following example questions were drawn from the MCQ synthetic benchmark constructed from arXiv_2502_17521v1 [45].

### v3

@@ Appendix A Expanded related works
When constructing knowledge or fact based evaluation benchmarks, there are two broad categories of answers: multiple choice and open ended. Open-ended questions provide the model less information and require synthesis (Balepur et al., 2025), but at the cost of needing another LM to evaluate the correctness of the results (LM-as-a-Judge (Xu et al., 2023)) (Myrzakhan et al., 2024). Multiple choice question (MCQ) results can be parsed and graded more simply. Modern evaluation frameworks like Inspect-AI (UK AI Security Institute, 2024) enable shuffling answer options and contain default prompts for both open-ended and MCQ evaluations. With MCQ, having the answers visible to the LM can artificially inflate performance (Robinson et al., 2022), even going as far as the model identifying the correct answers without seeing the question (Balepur et al., 2024). Both open-ended and multiple choice questions have weaknesses to question reformatting and phrasing where semantically identical questions produce different answers (Wang et al., 2024a). The simplest form of open-ended evaluation relies on a judge LM, though other frameworks have been proposed which attempt to improve the precision of open-ended evaluations (Bernard et al., 2024; Kamalloo et al., 2023).
Appendix B Replication of MMLU-pro
Because LM-as-a-judge was used throughout our experiments, a human evaluation of a subsampled set of questions generated to replicate MMLU-pro was performed in order to validate results. The authors took a random 20 questions from each evaluated subset of MMLU-pro (biology, chemistry, computer science, math, physics, and psychology) and manually graded the question/answer pair correctness and grounding against the source context from textbooks. Each question-answer pair was labeled with each of correct/incorrect and supported/unsupported by the context. Answer correctness was often negatively impacted by grounding failures, such that a question was missing information from the context required to be correctly answered. Per-domain details are as follows:
Physics: 19/20 questions correct, with the incorrect Question 11 incorrectly canceling the torque of a pivot. 14/20 questions were fully supported by the grounding context; of the remaining 6, the most common failure was the question-writing model relying on model parameter knowledge to fill in common knowledge which was not explicitly included in the grounding context. See Section˜B.1 for a per-question breakdown of the grounding failures in the physics category.
Biology: 20/20 question/answer pairs correct, and 20/20 with correct and valid grounding in the source contexts.
Chemistry: 18/20 question/answer pairs correct, 15/20 questions were fully grounded in the context. Both incorrect answers failed due to the question-writing model not including all required information in the question and or modifying minor details from the context. For example:
Question 13: Based on the descriptions provided, what are the chemical compositions of the protective passivation films that develop on the Statue of Liberty’s copper surface and on stainless steel, how does the text characterize their adherence to the underlying metal, and does it mention any self‑healing capability under cyclic wet‑dry exposure? Answer: The copper patina consists of copper oxides transformed by , , and into a blue‑green protective film, stainless steel’s passivation layer is chromium oxide, both are described as adherent protective layers, and the text does not address self‑healing under wet‑dry conditions. Human Evaluation: Incorrect and Unsupported because the question relies on information present in the context. Additionally, the answer formulation changed the context text from “The chromium tends to collect near the surface, where it corrodes and forms a passivating an oxide layer that protects the iron” to the wrong answer of “chromium oxide”.The five Chemistry grounding failures resulted from the question writing model added additional correct information from its parameter knowledge to flush out the question. For example:
Question 7: Using , predict how the spontaneity of NaCl melting changes when the temperature is raised from 500 °C to 700 °C, and explain why the entropy contribution becomes more decisive at the higher temperature. Answer: Nonspontaneous at 500 °C () and spontaneous at 700 °C () because the larger term at the higher temperature outweighs the positive . Human Evaluation: answer is Correct but Unsupported by the grounding context as its missing the equation. -
Computer science: 18/20 question/answer pairs correct, and 18/20 with correct and valid grounding in the source contexts. Both incorrect question/answer pairs were judged as failures because the question could not stand independent of the grounding context, as the question referenced information in the grounding context. These two grounding failures caused the questions to be labeled incorrect.
Math: 18/20 question/answer pairs correct, and 18/20 with correct and valid grounding in the source contexts. One of the two incorrect questions was the result of the question not restating required information from the context. The other incorrect question was more interesting, as the question-writing model wrote an answer that is open-ended and non-unique:
Question 7: Construct a rational equation that, after multiplying by its least common denominator, yields an algebraic solution which is exactly an excluded value, and explain why this equation has no real solution despite the apparent solution. Answer: ; the only solution is excluded because it zeroes the denominator, so there is no real solution. Human Evaluation: answer is Incorrect and Supported by the grounding context, because the provided answer is one possible answer of many.Both math subset grounding failures were a result of the question writing model adding additional correct information from its parameter knowledge to flesh out the question.
Psychology: 18/20 question/answer pairs correct, and 19/20 with correct and valid grounding in the source contexts.
Out of the 120 human evaluated question/answer pairs, 111 were correct. This gives the generative benchmarking an approximate noise floor. 104 of 120 questions had fully supported grounding in the document. The vast majority of the grounding failures fell into two categories: 1) the question writing model adding relevant correct information that was missing from the document, or 2) the generative question referencing the document and not being fully independent.
B.1 Detailed Human Evaluation Results: MMLU-pro Physics
In order to provide detailed insight into the questions and answers involved in the human evaluation, we reproduce the set of 20 manually evaluated physics questions here, along with the human judgment about question/answer quality.
Answer: Because water provides a low-resistance path that can complete a fault circuit, allowing current from the appliance to flow through a person in contact with the water.
Human Evaluation: Correct and Supported Question 2: How does the difference in attachment distance of the back muscles compared with the quadriceps alter their mechanical advantage during a lift, and why does using the quadriceps to raise the load markedly reduce the compressive force on the vertebrae as illustrated in the example?
Answer: Because back muscles have a short lever arm (low mechanical advantage) while quadriceps have a longer lever arm (high mechanical advantage), using the legs shifts the load to stronger joints and markedly lowers the force on the spine.
Human Evaluation: answer is Correct and Supported by the grounding context. Question 3: In a circuit where a switch is mistakenly wired in parallel with a resistive load, why does the load stay illuminated when the switch is open but go dark when the switch is closed, using Kirchhoff’s current law and the notion of a short circuit to justify the behavior?
Answer: With the switch open, KCL forces all current through the load giving it full voltage; closing the switch creates a short-circuit branch that shunts the current, leaving near-zero voltage across the load so it extinguishes.
Human Evaluation: answer is Correct and Supported by the grounding context. Question 4: For the high-pressure gas cylinder described (initial pressure , initial temperature ), determine (a) the pressure after isochoric cooling to dry-ice temperature , (b) the pressure if of the gas escapes, (c) the temperature to which the cylinder must be cooled to lower the pressure to without any gas loss, and (d) whether such cooling is a practical safety strategy.
Answer: After cooling ; after loss ; to reach , , a temperature far too low for practical use, making cooling impractical.
Human Evaluation: answer is Correct but Unsupported by the grounding context because its missing the governing relationships (ideal gas law / proportionalities). Question 5: Considering the decay processes of neutral and charged pions outlined, which component of an extensive air shower that spreads over several square kilometers retains the greatest share of the primary cosmic-ray energy, and why?
Answer: The electromagnetic component, because decays into photons that dominate the cascade.
Human Evaluation: answer is Correct but Unsupported by the grounding context because its missing the cosmic ray shower component energy shares. Question 6: Starting from the continuity of pressure and particle velocity at an interface between two media, derive the intensity reflection coefficient for an ultrasound wave and explain the physical significance of the numerator, denominator, and the squaring operation in the resulting expression.
Answer: ; and are the media’s acoustic impedances, reflects mismatch, represents total impedance at the interface, and squaring yields the intensity ratio.
Human Evaluation: answer is Correct and Supported by the grounding context. Question 7: In a mixed series-parallel network where a set of resistors is connected in parallel and that branch is in series with other elements, which tolerance extreme for the parallel resistors yields the highest total power dissipation, and why does this represent the worst-case overload condition?
Answer: When the parallel resistors are at their minimum tolerance (lowest resistance), because this minimizes equivalent resistance and maximizes total current and power.
Human Evaluation: answer is Correct and Supported by the grounding context. Question 8: During the inelastic collision of a moving disk with a stick pivoted at one end, why does the system’s angular momentum about the pivot remain conserved despite external forces, whereas its total linear momentum in the laboratory frame does not?
Answer: Because the external force passes through the pivot, giving zero torque (conserving angular momentum), but it supplies a net external impulse that changes linear momentum.
Human Evaluation: answer is Correct and Supported by the grounding context. Question 9: Why can a pion exchanged between a proton and a neutron transmit the strong nuclear force without being directly observable, and how does its fleeting existence satisfy while leaving the total energy–momentum of the nucleon system unchanged?
Answer: Because the pion exists only for a time (with ), allowing a temporary energy borrowing that is repaid when the pion is reabsorbed, so the nucleon system’s total energy–momentum stays conserved, rendering the pion virtual and unobservable.
Human Evaluation: answer is Correct and Supported by the grounding context. Question 10: When a cyclist applies a constant torque to the drivetrain, how does decreasing the wheels’ moment of inertia affect the bicycle’s linear acceleration, and what is the underlying reason?
Answer: Linear acceleration increases because a lower moment of inertia gives a higher angular acceleration for the same torque, reducing the rotational energy share and directing more power into translational motion.
Human Evaluation: answer is Correct and Supported by the grounding context. Question 11: When bending forward to touch the toes while keeping the heels and back against a wall, what torque must the hip and knee muscles generate about the heels to keep the system in static equilibrium, and how does selecting the wall contact point as the pivot affect this torque calculation?
Answer: A torque equal to the body’s weight multiplied by its horizontal distance from the heels; choosing the wall contact as pivot removes this torque because the wall reaction provides an equal balancing moment.
Human Evaluation: answer is Incorrect but the question is Supported by the grounding context. Question 12: Derive, using the subshell capacity formula , why the total number of electrons that can occupy all subshells of a shell with principal quantum number equals , and list the implicit assumptions required for this derivation.
Answer: Summing for to yields ; the derivation assumes the allowed -range (), full -degeneracy (), spin degeneracy of 2, independent full filling of subshells, and neglects relativistic and electron-interaction effects.
Human Evaluation: answer is Correct and Supported by the grounding context. Question 13: Using , determine how the relativistic kinetic energy scales as approaches , and state the limiting value of the classical kinetic energy in the same limit.
Answer: Relativistic diverges to infinity, while classical approaches (a finite value).
Human Evaluation: answer is Correct and Unsupported by the grounding context as its missing information about the classical limiting value. Question 14: Which term describes chaotic systems that nevertheless develop organized, long-lasting structures, and what is the example of such a structure in Jupiter’s atmosphere?
Answer: Stable chaos, exemplified by the Great Red Spot on Jupiter.
Human Evaluation: answer is Correct and Supported by the grounding context. Question 15: Derive the decibel expression for sound-intensity level from the fact that human hearing responds logarithmically to intensity, and explain why the formula uses a factor of 10 for intensity ratios but a factor of 20 when expressed in terms of pressure ratios.
Answer: (the factor 20 arises because intensity pressure2).
Human Evaluation: answer is Correct and Supported by the grounding context. Question 16: Which analytical method, involving the decomposition of the initial velocity into horizontal and vertical components, is used to determine a projectile’s range, and what explicit assumption about air resistance does the context provide for this calculation?
Answer: The analytical method that decomposes the initial velocity into horizontal and vertical components, with air resistance assumed zero.
Human Evaluation: answer is Correct and Supported by the grounding context but the question is degenerate and directly references the context. Question 17: How can one determine the fraction of a human body that remains submerged when floating in freshwater versus seawater based solely on the body’s overall density, and why does an increase in low-density fat relative to higher-density muscle cause that fraction to decrease (i.e., make the person float higher)?
Answer: The submerged fraction equals the body’s density divided by the fluid’s density ( in fresh water and in salt water); increasing low-density fat lowers , thus decreasing that fraction and raising the body higher in the water.
Human Evaluation: answer is Correct but Unsupported by the grounding context because its missing the density fraction formula. Question 18: How does defining internal energy microscopically as the total kinetic and potential energy of a system’s atoms and molecules justify the macroscopic first-law relation , and why does this make path-independent while heat and work remain path-dependent?
Answer: Because internal energy is a state function determined solely by the microscopic kinetic and potential energies of the particles, its change equals the net heat added minus the net work done () for any process, making path-independent while and are path-dependent.
Human Evaluation: answer is Correct and Supported by the grounding context. Question 19: Considering Germany’s 2020 renewable electricity share of and the projected rise in global energy demand, what does achieving the 2030 target of renewable power imply about its technical and economic feasibility?
Answer: It implies an extremely challenging technical and economic undertaking, requiring about a 27-fold increase in renewable share.
Human Evaluation: answer is Correct and Supported by the grounding context but the context table is missing some renewable information which makes the answer correct only w.r.t. the context. Additionally, the question is degenerate and uninteresting. Question 20: A pacemaker lead moving at perpendicular to an MRI scanner’s magnetic field generates a Hall voltage of . What magnetic field strength does this imply, and, accounting for the lead’s inductance and the rapid switching of MRI gradient fields, is the induced current likely to disrupt pacemaker operation?
Answer: , and the induced current is negligible, so pacemaker interference is unlikely.
Human Evaluation: answer is Correct but Unsupported by the grounding context because its missing information about lead inductance and MRI switching.
Appendix C Impact of question and answer length
@@
While longer more detailed questions (both MC and OE) produce a model over-performance (see Figure˜6), the same trend does not appear in answer length. Figure˜9 demonstrates this via scatterplot showing synthetic benchmark accuracy over-performance. Only the gpt-4.1-mini at the very right of the plot is significantly biased above the expected trend line. The over-performance trend shown for questions in Figure˜6(b) is absent.
@@
The model gpt-4.1-mini demonstrates the largest over-performance for both question length and answer length as shown in Figure˜6(b). The following question answer pair was generated by gpt-4.1-mini (37 tokens) using a context from Squadv2, highlighting the difference in model verbosity in both question and answers:
@@
Appendix D Synthetic benchmark question diversity
@@
Most LMs produce similar diversity scores that are closely aligned with the human questions Some highly capable models (such as gpt-4.1) show comparatively worse question diversity while performing well in clarity, correctness, and groundedness. Figure˜10 shows the question diversity per generating model along with the original question diversity.
Appendix E Multiple-choice vs open-ended questions
@@
The generating models which create overly easy questions by including too much detail (like phi-4 and gpt-4.1-mini) do so for both open ended and multiple choice questions. The long question problem becomes even more pronounced for open ended questions as shown in Figure˜11.
@@
These overly long synthetic questions result in an over-performance by smaller models on the generated benchmark as shown in Figure˜12. We hypothesize this happens because the additional detail in the question provides the LM under evaluation more information to pattern match against text seen in the training phase, thereby increasing the likelihood that the LM retrieves memorized patterns relevant to the question. This effect is seen in Figure˜12(a) for multiple choice, and Figure˜12(b) for open-ended. Both plots in Figure˜12 demonstrate LM benchmark evaluation accuracy over-performance ().
@@
While similar trends in question length and the over-performance that induces, the evaluation accuracy correlations and model rank order work equally well for open ended and multiple choice questions. Figure˜13(a) plots the correlation in evaluation accuracy between the human reformatted questions and the synthetic generated benchmark for multiple choice, and Figure˜13(b) for open-ended.
Appendix F Evaluation benchmark accuracy and ranking correlation
@@
Figure˜13(a) shows the evaluation accuracy scatterplot comparing the reformatted human questions against the synthetic generated benchmark questions for the bolded ensemble models in Table˜2. This ensemble produces a Pearson accuracy correlation of between the evaluation accuracies for synthetic and human benchmarks. Each plot marker is the average evaluation accuracy for a given model-dataset combination. While this model ensemble produces reasonable overall accuracy correlation, occasionally specific LMs generate undesirable benchmarks. Figure˜14 showcases a failure where the synthetic benchmark generated by gpt-4.1-mini contains overly easy questions where all models perform near evaluation accuracy (Pearson accuracy correlation .). This over-performance is likely linked to the question length problem.
@@
Figure˜15 plots the evaluation model performance across all datasets, comparing the synthetic and human accuracy Spearman ranking correlation of the evaluation model capability. This matches the use case of a user selecting between various models for a specific application relevant to the provided grounding documents. Figure˜15(a) is the rank order barchart equivalent of Figure˜13(a) and Figure˜15(b) is the rank order barchart equivalent of the failure case in Figure˜14.
@@
Table˜4 showcases the Spearman rank correlation between evaluation benchmarks written by humans and the synthetic data versions. The table includes data for all three of the human question reformatting models used to demonstrate the minimal impact that reformatting has on the overall high level results. The reformatting operation was evaluated against a subset of datasets including Squadv2, HotpotQA, TrivaQA-web, NaturalQuestionsShort, PubMedQA, and SecQA. Overall, the best outcome is to ensemble together a few questions generated by various reasonably capable LMs to avoid the known slight self-bias/self-preference issue. While we used Llama-4-Maverick as the text feature extractor for measures like groundedness and clarity, that model was unexpectedly dominant in these final results.
@@
Table˜5 contains the benchmark evaluation Pearson accuracy correlation between the reformatted human questions and the synthetic data generated questions. Similarly to the Spearman ranking, Llama-4-Maverick performed very well in these tests. However, unlike the ranking data, using a ensemble of models produces significantly worse Pearson accuracy correlation numbers under some circumstances. Likely due to the long and overly easy questions written by gpt-4.1 and o4-mini.
@@
Appendix G Question clarity and groundedness
@@
Appendix H Case study on academic papers
Our approach was demonstrated on two academic papers: “Current Solutions and Future Trends for Robotic Prosthetic Hands” (labeled annurev-control-071020-104336) (Mendez et al., 2021), “Recent Advances in Large Language Model Benchmarks against Data Contamination: From Static to Dynamic Evaluation” (labeled arXiv_2502_17521v1) (Chen et al., 2025) and the recently released “America’s AI Action Plan” (4). Figure˜16 showcases LM benchmark evaluation performance on these two datasets; the MCQ results show most models performing well before a sharp dropoff with 4B and fewer parameter models and the Open-Ended results with a much smoother decline in performance. Multiple-choice questions can be easier to answer due to the presence of the answer options, even if the evaluation framework shuffles the option order. This effect is likely the cause of the MCQ evaluation accuracy to remaining higher for longer.
The following example questions were drawn from the MCQ synthetic benchmark constructed from arXiv_2502_17521v1 (Chen et al., 2025).